Can You please help me this these problems
1. Solve the equation using addition (elimination) method. If the answere is an unique solution present it as an ordered pair:(x,y),if not specify whether the answere is "no solution" or "infinitely many solution"
5x-4y =1
- 10x +8y + =-3
2. Solve the equation by using the substitution method. If the answere is a unique solution present it as an ordered pair. If not specify whether it "no solution" or "infinitely many solution.
7x + y = 1
3x + 2y = -9

Answer by oscargut(2103)

1. Solve the equation using addition (elimination) method. If the answere is an unique solution present it as an ordered pair:(x,y),if not specify whether the answere is "no solution" or "infinitely many solution"
5x-4y =1
- 10x +8y + =-3
then doing 1st eq by 2
10x-8y=2
-10x+8y=-3
adding both equations
0=-1 then the system has "no solution"
2. Solve the equation by using the substitution method. If the answere is a unique solution present it as an ordered pair. If not specify whether it "no solution" or "infinitely many solution.
7x + y = 1
3x + 2y = -9
from the 1st eq we have y=-7x+1
then using 2nd eq we have 3x + 2(-7x+1)=-9
-11x+2=-9
-11x=-11
x=1
and y=-7(1)+1=-6
Then solution is (1,-6)